OpenStudy (anonymous):

p − 4 = −9 + p Could someone help me solve this two-step equation? Thank you.

OpenStudy (anonymous):

Since the variables on both sides are same therefore this problem won't have any solution.

OpenStudy (anonymous):

I know. But I have to explain how to solve it & in my own words how there is no solution.

OpenStudy (anonymous):

Ok pls wait

OpenStudy (anonymous):

Alright. Thank you.

OpenStudy (anonymous):

p − 4 = −9 + p i.e. p-p=4-9 \[i.e. 0\neq-5\]

OpenStudy (anonymous):

How in words is there not a solution?

OpenStudy (anonymous):

@dpasingh

hero (hero):

I wish you would have stayed at the previous question for two seconds longer. I wasted my time typing a response to your question. http://openstudy.com/study#/updates/51cd80fee4b063c0de5b3dc2